# GraphPilot: Grounded Scene Graph Conditioning for Language-Based Autonomous Driving

### v2

@@ 3.2 Scene Graph Serialization
Figure 3: Serialization formats. We serialize scene graphs as Text, JSON, or YAML, each encoding subject-predicate-object triplets. Text uses compact natural-language packing for brevity, JSON provides a structured and parser-friendly representation, and YAML achieves lower token counts through minimal syntax.
@@
JSON. The JSON form mirrors the graph with two arrays: nodes and links. Each node stores a stable identifier and semantic class (id, base_class). Each link stores source, target, and a labels array containing all predicates for that ordered pair. JSON is therefore packed at the pair level with one link per $(s,o)$ with possibly multiple labels, so it does not apply membership grouping. The result is a uniform, parser-friendly schema.
YAML. The YAML form is an indentation-based analogue of the JSON schema with the same two top-level collections (nodes, links) and the same fields. Like JSON, it represents one ordered pair per link and aggregates predicates in a labels list. Owing to indentation rather than braces and fewer quotation marks, YAML typically yields fewer tokens than JSON while preserving identical semantics.
@@ 3.3 Scene Graph Injection
We instantiate three prompt templates that differ in layout and instruction framing while keeping the underlying content identical. Fig. 4 shows all three templates side by side for direct comparison.
The first template (V1) concatenates the navigation command and the scene graph in a single line without section headers. The command appears first, followed by a short cue and the serialized graph. This variant minimizes prompt overhead and matches instruction plus context compositions seen in prior work. The second template (V2) introduces explicit section headers for the scene graph and for the navigation command and begins with a reminder that the model acts as the ego vehicle. The graph is shown as a headed block and the command as a separate headed block, which makes both inputs distinct and separable when contexts become longer. The third template (V3) keeps the sectioning and wraps the scene graph in a fenced block with a format prefix such as text, json, or yaml. It also states a short Primary Objective that emphasizes adherence to the external navigation command. The fence treats the graph as read only and aims to stabilize tokenization for larger graphs or mixed formatting, and the format prefix encourages consistent parsing across representations.
@@ 3.4 Scene Graph Conditioning
We condition the model on structured relational information by including a serialized scene graph alongside the natural-language navigation command in the input prompt. In this work, we use the term conditioning to refer specifically to prompt-level supervision: the model receives explicit relational context as part of its input, but its architecture, training objective, and loss functions remain unchanged. This prompt-based conditioning allows the model to ground decision making and planning in spatial topology, regulatory context, and inter-actor interactions that would otherwise need to be inferred implicitly from visual features alone. Since scene graph-conditioning operates purely at the input level, it is model-agnostic and does not require modifying perception modules, introducing graph encoders, or defining auxiliary training losses.
@@ 4 Experiments
We evaluate how scene graph-conditioning affects the driving performance of VLM-based autonomous driving systems across multiple configurations and training setups.
@@ 4.1 Benchmark
We evaluate on the public LangAuto benchmark [shao2024lmdrive], which adapts the CARLA Leaderboard for language-based driving. The benchmark covers eight towns with diverse traffic layouts, weather, and illumination, and replaces discrete high-level commands with natural-language navigation instructions. Routes are partitioned by length into Tiny (<150m<150m), Short (150500m150-500m), and Long (>500m>500m) tracks.
Metrics. We follow the official CARLA Leaderboard metrics. Route completion (RC) measures the fraction of route distance achieved. Infraction score (IS) applies multiplicative penalties for collisions, traffic-rule violations, excessive route deviation, and timeouts, starting from 1.0 and decaying with event type and count. The driving score (DS) for a route is the product of RC and IS. We report the mean across routes as the overall benchmark score. Following the LangAuto protocol, each configuration is executed three times and results are averaged.
@@ 4.2 Baselines
We use LMDrive [shao2024lmdrive] and BEVDriver [winter2025bevdriver] as baselines. Both systems represent state-of-the-art VLM-based autonomous driving approaches and are designed to process both visual and natural language inputs. In our experiments, we use the officially released model variants: LMDrive with LLaVA-v1.5 [liu2024llava] and BEVDriver with LLaMA-7B [touvron2023llama].
Dataset Collection. The publicly available LMDrive dataset does not include scene graphs. Therefore, we extend the official LMDrive data generation pipeline with our scene graph generation, abstraction, and serialization modules (Sec. 3) so that ground-truth scene graphs are captured alongside the original modalities. The resulting data preserves the original distribution over selected routes, towns, weather, and illumination, and remains fully compatible with both LMDrive and BEVDriver.
@@ 4.3 System Configurations
For our approach, we define four scene graph usage settings, i.e., SG00, SG01, SG10, SG11, to evaluate how relational supervision affects language-based driving.
SG00 uses no scene graphs during training or test-time and serves as our main baseline. We distinguish between two variants: SG00-PT uses the original pretrained checkpoints provided by the model authors, while SG00-FT further fine-tunes these checkpoints on our collected dataset. This distinction ensures a fair comparison, as directly comparing our scene graph-conditioned models to SG00-PT would conflate the effects of additional training data with those of scene graph supervision. SG01 includes scene graphs only at test-time. This setting probes whether pretrained models can benefit from relational structure without any prior exposure to scene graphs during training. SG10 includes scene graphs during training but omits them at test-time. This measures whether models can internalize relational priors and generalize without needing explicit scene graph input during inference. SG11 includes scene graphs both during training and test-time and represents the full supervision scenario.
@@ 4.4 Implementation Details
We initialize all models from the officially released checkpoints. The language backbone is fine-tuned using LoRA adapters, while the vision encoder remains frozen. Preliminary experiments revealed that jointly optimizing both backbones or including auxiliary notice instructions during training as used in some baseline variants led to reduced closed-loop driving performance. Therefore, we omit notice instructions entirely and keep the vision encoder fixed throughout. Accordingly, we follow the default optimization settings of each baseline and do not introduce additional training losses or hyperparameter tuning strategies. All components outside the prompt remain unchanged.
@@ 4.5 Training
All models are trained on eight NVIDIA L40s GPUs. We use the same train/validation split as in the original works and train for five epochs with a global batch size of eight.
When a scene graph is used during training (SG11 or SG10), it is integrated based on the selected abstraction level, serialization format, and injection template. These choices govern the form and placement of the graph in the prompt but do not alter the model architecture or optimization procedure. The total training time (GPU hours) per run depends primarily on the abstraction level and serialization format, ranging from 61 hours (Actor-Only, Text) to 107 hours (Full, JSON) for LMDrive. Among the formats, Text is the most efficient, followed by YAML, while JSON incurs the highest overhead due to its more verbose structure. Injection template has negligible effect on training time.
4.6 Results
Table 1: Baseline performance. No scene graphs are used during training or test-time. SG00-PT refers to the official pretrained checkpoints, while SG00-FT applies further fine-tuning on our collected dataset. LMDrive degrades under SG00 fine-tuning, whereas BEVDriver improves, highlighting model-specific sensitivity to fine-tuning without relational supervision. Best model-specific results are in bold.
Method DS $\uparrow$ RC $\uparrow$ IS $\uparrow$
LMDrive (SG00-PT) 42.95 54.81 0.79
LMDrive (SG00-FT) 41.55 48.22 0.88
BEVDriver (SG00-PT) 44.70 49.70 0.90
BEVDriver (SG00-FT) 48.94 56.37 0.86
Baselines. We first establish baselines by using the official implementations and pretrained checkpoints of LMDrive and BEVDriver, referred to as SG00-PT. These results serve as primary reference points to ensure fair, environment-matched comparisons. [^1] We report reproduced results rather than official numbers, as the authors of LMDrive acknowledge that performance varies significantly depending on GPU model as reported by other authors as well (see issues 37, 52, and 56 in LMDrive repository; links omitted for review according to author guidelines). Same behaviour was observed for BEVDriver. Additionally, to isolate the effect of scene graph supervision itself, we fine-tune both models on our dataset without scene graph input at training or test-time (SG00-FT). Tab. 1 shows that SG00 fine-tuning shifts performance in a model-specific way: LMDrive gains in safety (higher IS) but sees a drop in RC, while BEVDriver substantially improves RC and resulting in a higher DS. These outcomes provide a neutral reference for assessing the added value of scene graph conditioning.
Table 2: Test-time-only scene graph injection (SG01). LMDrive performance using different combinations of abstraction levels, serialization formats, and injection templates. Results averaged across LangAuto Tiny, Short, and Long. Actor-Only scene graphs paired with the V3 injection template consistently perform best across formats, demonstrating that concise but structured relational input yields driving improvements even without additional training. Best mean results are highlighted in bold.
Serialization Abstraction V1 V2 V3 Mean
@@ 4.6 Results
Text Full 38.9 46.2 0.84 41.3 49.7 0.83 41.8 49.4 0.83 40.7 48.4 0.83
Road-Level 40.6 47.0 0.86 40.5 50.1 0.81 41.5 49.4 0.83 40.9 48.8 0.83
Actor-Only 40.0 48.7 0.85 43.1 51.2 0.84 44.7 53.0 0.84 42.6 51.0 0.84
JSON Full 34.6 42.2 0.84 41.5 48.7 0.84 44.6 54.6 0.83 40.2 48.5 0.84
Road-Level 34.4 43.4 0.83 41.8 51.2 0.81 46.6 56.3 0.82 41.0 50.3 0.82
Actor-Only 40.3 47.3 0.85 42.6 52.3 0.82 44.7 55.6 0.82 42.6 51.7 0.83
YAML Full 39.4 45.5 0.85 41.8 50.4 0.82 41.8 51.9 0.80 41.0 49.3 0.82
Road-Level 39.2 46.6 0.85 41.8 51.5 0.82 42.1 50.6 0.83 41.1 49.6 0.83
Actor-Only 42.8 52.0 0.84 43.6 53.2 0.82 44.7 52.7 0.84 43.7 52.6 0.83
@@
Actor-Only 41.1 49.4 0.85 43.1 52.2 0.82 44.7 53.8 0.83 43.0 51.8 0.84
Table 3: Token statistics for scene graph variants. Mean and maximum token counts for each abstraction level and serialization format in the training set, measured using the LLaVA-v1.5 tokenizer, showing that abstraction level and serialization format substantially affect token count.
Abstraction Text JSON YAML
Mean Max Mean Max Mean Max
Full 437 1605 2905 10020 1712 5950
Road-Level 370 1576 2172 8753 1246 5052
Actor-Only 69 607 409 2591 242 1606
@@
Test-Time-Only Scene Graph Injection (SG01). We begin by evaluating which scene graph integration strategies are most effective when applied at test-time only without additional training (SG01), to assess whether a pretrained model can benefit from relational context without having been explicitly conditioned on it during training.
@@
Specifically, we evaluate all combinations of abstraction level, serialization format, and injection template. Tab. 2 reports DS, RC, and IS for LMDrive averaged over the LangAuto Tiny, Short, and Long tracks. We give mean rows and columns to support comparisons across serialization formats and injection templates, respectively. Notably, multiple test-time scene graph configurations already surpass the reproduced baseline without scene graphs (Tab. 1), underscoring the potential of this approach even without additional training. Two consistent trends emerge: First, injection template V3 outperforms both V2 and V1, with V1 trailing significantly. Second, Actor-Only abstractions yield the best results across all formats, indicating that minimal yet structured relational information is most beneficial when applied to pretrained models.
@@
These results highlight the strength of lean representations and well-designed prompts, even without additional training. Actor-Only graphs in particular strike a favorable balance between semantic relevance and token budget, introducing the fewest additional tokens across all formats, as shown in Tab. 3. Based on these insights, we select the Actor-Only abstraction and injection templates V2 and V3 for all subsequent fine-tuning experiments, focusing on configurations that maximize prompt efficiency and clarity.
Table 4: Scene graph-conditioned fine-tuning. Comparison of SG10 and SG11 using Actor-Only graphs with V2 and V3 prompt templates. SG10 matches or exceeds SG11 in most settings, showing that models retain relational grounding and can drive effectively without test-time scene graphs. Best model-specific results are highlighted in bold.
Serialization Configuration LMDrive BEVDriver

### v3

@@ 3.2 Scene Graph Serialization
Figure 3: Serialization formats. We serialize scene graphs as Text, JSON, or YAML, each encoding subject-predicate-object triplets.
@@
Structured Formats (JSON/YAML). Both formats mirror the graph using two collections: nodes (storing id and base_class) and links. To ensure a compact schema, each link represents a unique ordered pair $(s,o)$ and aggregates all applicable predicates into a single labels list. While JSON provides a standard parser-friendly structure, YAML uses indentation-based nesting, which yields a more token-efficient representation by omitting braces and quotes.
@@ 3.3 Scene Graph Injection
We instantiate three prompt templates that differ in layout and instruction framing while keeping the underlying content identical. Fig. 4 shows all three templates side by side for direct comparison. The first template (V1) concatenates the navigation command and the scene graph in a single line without section headers. The command appears first, followed by a short cue and the serialized graph. This variant minimizes prompt overhead and matches the instruction-plus-context compositions observed in prior work. The second template (V2) introduces explicit section headers for the scene graph and for the navigation command and begins with a reminder that the model acts as the ego vehicle. The graph is shown as a headed block and the command as a separate headed block, which makes both inputs distinct and separable when contexts become longer. The third template (V3) keeps the sectioning and wraps the scene graph in a fenced block with a format prefix such as text, json, or yaml. It also states a short Primary Objective that emphasizes adherence to the external navigation command. The fence treats the graph as read-only and aims to stabilize tokenization for larger graphs or mixed formatting, and the format prefix encourages consistent parsing across representations.
@@ 3.4 Scene Graph Conditioning
We condition the model on structured relational information by including a serialized scene graph alongside the natural-language navigation command in the input prompt. In this work, we use the term conditioning to refer specifically to prompt-level supervision: the model receives explicit relational context as part of its input, but its architecture, training objective, and loss functions remain unchanged. This prompt-based conditioning allows the model to ground decision-making and planning in spatial topology, regulatory context, and inter-actor interactions that would otherwise need to be inferred implicitly from visual features alone. Since scene graph-conditioning operates purely at the input level, it is model-agnostic and does not require modifying perception modules, introducing graph encoders, or defining auxiliary training losses.
@@ 4 Experiments
We evaluate the impact of scene graph-conditioning on VLM-based autonomous driving systems across multiple configurations and training setups.
@@ 4.1 Benchmark
We evaluate our approach on two CARLA-based [6] closed-loop benchmarks: LangAuto [36] and Bench2Drive [17]. LangAuto adapts the CARLA Leaderboard for language-based driving across eight towns, featuring diverse traffic, weather, and illumination conditions, in which discrete goal points are replaced by natural-language instructions. Bench2Drive utilizes 220 short routes, approximately 150m each, spanning 44 interactive scenarios with five routes per scenario across 15 towns to evaluate state-of-the-art driving agents.
Metrics. For both benchmarks, we report the official Driving Score (DS), defined as the product of Route Completion (RC) and the Infraction Score (IS). RC measures the fraction of the route distance achieved, while IS applies multiplicative penalties for traffic-rule violations and collisions. We also report the Success Rate (SR) for Bench2Drive, which measures the proportion of routes successfully completed. Following standard protocols, all results are averaged across three execution seeds.
@@ 4.2 Baselines
We evaluate our approach using three state-of-the-art VLM-based driving systems: LMDrive [36] (LLaVA-v1.5 [23]) and BEVDriver [48] (LLaMA-7B [43]) for LangAuto, and SimLingo [33] (InternVL2-1B [8]) for Bench2Drive. These baselines represent diverse architectural paradigms, allowing us to demonstrate the broad applicability of our method across different vision-language foundations. Following standard protocols, all baselines are evaluated using their officially released model variants to ensure a fair comparison.
Dataset Collection. The publicly available LMDrive and SimLingo datasets do not include scene graphs. Therefore, we extend the official data generation pipelines with our scene graph generation, abstraction, and serialization modules so that ground-truth scene graphs are captured alongside the original modalities. These graphs are constructed from privileged simulator state by utilizing 3D bounding boxes and HD map metadata to ensure high-fidelity relational labels. While our pipeline is implemented for CARLA, this approach remains directly applicable to real-world datasets (e.g., nuScenes) where such annotations are standard. The resulting data preserves the original distribution over selected routes, towns, weather, and illumination, and remains fully compatible with the original frameworks. We utilize our extended LMDrive dataset to train LMDrive and BEVDriver, while the extended SimLingo dataset is used to train SimLingo.
@@ 4.3 System Configurations
For our approach, we define four scene graph usage settings, i.e., SG00, SG01, SG10, and SG11, to evaluate the impact of relational supervision on driving. SG00 uses no scene graphs during training or test-time and serves as our main baseline. We distinguish between two variants: SG00-PT uses the original pretrained checkpoints provided by the model authors, while SG00-FT further fine-tunes these checkpoints on our collected dataset. This distinction ensures a fair comparison, as directly comparing our scene graph-conditioned models to SG00-PT would conflate the effects of additional training data with those of scene graph supervision. SG01 includes scene graphs only at test-time. This setting probes whether pretrained models can benefit from relational structure without any prior exposure to scene graphs during training. SG10 includes scene graphs during training but omits them at test-time. This measures whether models can internalize relational priors and generalize without needing explicit scene graph input during inference. SG11 includes scene graphs at both training and test-time and represents the fully supervised scenario.
@@ 4.4 Implementation Details
We initialize all models from officially released checkpoints and strictly adhere to the optimization protocols provided by each baseline. While all models train their respective projection layers and action heads, they differ in their primary backbones: LMDrive keeps both the vision encoder and language model frozen while training the Q-Former [21]; BEVDriver freezes the vision encoder but trains the LLM via LoRA [14] and the Q-Former; and SimLingo utilizes a fully trainable vision encoder alongside a LoRA-optimized language backbone. We maintain these original training recipes without additional losses or auxiliary tasks. By ensuring that all components outside the prompt remain unchanged, performance gains are strictly attributable to the internalized relational knowledge of our scene graphs, rather than to architectural modifications.
@@ 4.5 Training
All models are trained on eight NVIDIA L40s GPUs for five epochs using the original train/validation splits. Scene graphs (SG10/SG11) are integrated via the prompt based on the selected abstraction and serialization without altering the underlying model architecture. Total training time depends primarily on the baseline architecture and serialization format, with Text being significantly more efficient than the verbose JSON format. Specifically, training ranges from 20h (AO, Text) to 29h (AO, JSON) for SimLingo. For BEVDriver, durations range from 34h (AO, Text) to 78h (Full, JSON), while LMDrive requires 61h (AO, Text) to 106h (Full, JSON). The choice of injection template has a negligible effect on training efficiency.
4.6 Results and Ablations
Table 1: Zero-shot test-time scene graph injection (SG01). Results for the pretrained LMDrive model across abstraction levels, formats, and templates. Results are averaged across LangAuto Tiny, Short, and Long. The rightmost column shows the mean token count for each configuration on the LMDrive dataset.
Serial. Abstract. V1 V2 V3 Mean Tokens
@@ 4.6 Results and Ablations
Text Full 38.9 46.2 0.84 41.3 49.7 0.83 41.8 49.4 0.83 40.7 48.4 0.83 437
Road-Level 40.6 47.0 0.86 40.5 50.1 0.81 41.5 49.4 0.83 40.9 48.8 0.83 370
[…]
JSON Full 34.6 42.2 0.84 41.5 48.7 0.84 44.6 54.6 0.83 40.2 48.5 0.84 2905
Road-Level 34.4 43.4 0.83 41.8 51.2 0.81 46.6 56.3 0.82 41.0 50.3 0.82 2172
[…]
YAML Full 39.4 45.5 0.85 41.8 50.4 0.82 41.8 51.9 0.80 41.0 49.3 0.82 1712
Road-Level 39.2 46.6 0.85 41.8 51.5 0.82 42.1 50.6 0.83 41.1 49.6 0.83 1246
[…]
@@
[…]
@@
Zero-Shot Test-Time Scene Graph Injection (SG01). We begin by evaluating which scene graph integration strategies are most effective when applied at test-time only, without additional training, to assess whether a pretrained model can benefit from relational context without prior explicit conditioning. This analysis identifies the optimal configuration for subsequent experiments.
@@
Specifically, we evaluate all combinations of abstraction level, serialization format, and injection template. Tab. 1 reports DS, RC, and IS for LMDrive averaged over the LangAuto Tiny, Short, and Long tracks. We provide mean rows and columns to facilitate comparisons across serialization formats and injection templates. Notably, multiple test-time configurations already surpass the performance of the original pretrained baselines, underscoring the inherent potential of relational grounding even without fine-tuning.
Figure 5: Scene graph abstractions. Comparison of Full, Road-Level, and Actor-Only abstractions, demonstrating how structural detail is progressively filtered.
@@
Two consistent trends emerge from this zero-shot evaluation. First, injection template V3 consistently outperforms V2 and V1, with V1 trailing significantly. Second, Actor-Only abstractions yield the best results across all formats, indicating that minimal but structured relational information is most beneficial for pretrained models. These findings highlight the strength of lean representations. Actor-Only graphs strike a favorable balance between semantic relevance and token budget by introducing the fewest additional tokens, as depicted in Fig. 5. Based on these insights, we select the Actor-Only abstraction and injection templates V2 and V3 for all subsequent fine-tuning experiments, focusing on configurations that maximize prompt efficiency and clarity.
Knowledge Internalization (SG10). Tab. 2 reports the driving performance for our SG10 configurations using the optimal Actor-Only abstraction and injection templates (V2, V3). To ensure fair, environment-matched comparisons, we reproduce all results on our local hardware using the original pretrained baselines (SG00-PT) and models fine-tuned without scene graphs (SG00-FT). This setup allows us to assess the extent to which structured relational context is internalized during training to benefit test-time, where graphs are absent.
Table 2: Knowledge internalization (SG10). Performance of our SG10 strategy across LMDrive, BEVDriver, and SimLingo using Actor-Only graphs compared to pretrained (SG00-PT) and fine-tuned (SG00-FT) baselines.
Config. LMDrive BEVDriver SimLingo
V2 V3 V2 V3 V2 V3
DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ SR $\uparrow$ DS $\uparrow$ SR $\uparrow$
SG00-PT 42.9 54.8 0.80 42.9 54.8 0.80 44.7 49.7 0.90 44.7 49.7 0.90 87.4 0.72 87.4 0.72
SG00-FT 41.5 48.2 0.88 41.5 48.2 0.88 48.9 56.4 0.86 48.9 56.4 0.86 87.4 0.71 87.4 0.71
Text 46.8 56.2 0.84 51.8 59.4 0.86 51.7 59.8 0.87 51.5 60.0 0.86 88.4 0.72 88.1 0.71
JSON 45.7 53.5 0.85 45.5 54.0 0.85 47.3 52.1 0.90 42.2 50.7 0.85 88.9 0.73 88.4 0.72
YAML 47.9 57.4 0.83 46.8 54.8 0.84 56.1 64.1 0.87 48.2 55.7 0.85 88.2 0.71 87.9 0.70
[…]
For LMDrive, our SG10 approach consistently outperforms both the pretrained baseline and the standard fine-tuning setup. The best configuration (Text, V3) achieves a DS of 51.8, marking a substantial improvement over the baseline (42.9) and the SG00-FT variant (41.5). Notably, these results significantly surpass the best zero-shot SG01 performance (44.7) previously shown in Tab. 1, demonstrating that explicit training with relational supervision is more effective than simple test-time injection. BEVDriver shows a similar trend of significant improvement. The top configuration (YAML, V2) reaches a DS of 56.1, outperforming its respective SG00-PT (44.7) and SG00-FT (48.9) benchmarks. While BEVDriver shows a preference for YAML serialization and the V2 template, the underlying pattern remains consistent with LMDrive: training-time scene graph exposure enhances internal planning behavior even when that structure is removed during deployment. To demonstrate that these benefits are not specific to the LangAuto benchmark, we further evaluate our strategy using SimLingo on the Bench2Drive dataset. Despite a high baseline performance of 87.4, SimLingo achieves consistent improvements across all formats, with the best configuration (JSON, V2) reaching a DS of 88.9. We hypothesize that scene graph supervision acts as a relational bottleneck during training. By forcing the model to predict actions conditioned on explicit subject-predicate-object triplets, the vision-language backbone (e.g., the Q-Former or LoRA-tuned LLM) learns to prioritize safety-critical spatial features and dynamic inter-actor dependencies. Consequently, the model ”learns to attend” to these relational cues within the raw sensory input, allowing the internalized reasoning logic to persist even when the text-based scene graph is removed at test-time. This SG10 strategy offers major practical advantages: it eliminates complex and potentially error-prone test-time scene graph generation and reduces prompt length, significantly lowering the computational and memory load essential for real-time systems.
Table 3: Extended infraction analysis. Comparison of fine-tuned LMDrive models on safety and behavior sub-metrics, averaged over V3-based configurations.
SG00-PT 42.94 54.81 0.80 0.08 2.83 4.28 2.31 0.00 3.79 11.95 2.93 1.63
SG00-FT 41.54 48.22 0.88 0.00 0.85 2.93 0.10 0.00 2.86 9.94 4.31 2.51
[…]
CP: Collision Pedestrians, CV: Collision Vehicles, CL: Collision Layout, RL: Red Light, SS: Stop Sign, Off: Off-Road, RD: Route Deviation, TO: Timeout, AB: Agent Blocked.
Figure 6: Qualitative behavior comparison. In each scenario, the LMDrive baseline (SG00-PT) on the left is compared against our SG10 on the right. The baseline predicts an unsafe lane change into traffic while failing to anticipate a right turn (Left Pair), and incorrectly turns right instead of left, whereas subsequently entering the opposite lane (Right Pair). In both scenarios, SG10 demonstrates superior command grounding and relational awareness.
Extended Infraction Analysis. To better understand how scene graph conditioning influences safety, we analyze detailed sub-metrics in Tab. 3, averaging over all V3-based LMDrive configurations. Compared to the original baseline (SG00-PT), SG10 yields significant reductions in critical infractions: vehicle (CV) and layout (CL) collisions drop by over 50%, while pedestrians (CP) collisions are nearly eliminated, see Fig. 6 left. Route Deviation (RD) also improves from 11.95 to 8.22, suggesting that training-time scene graph exposure fosters more structured driving even when graphs are absent at test-time. Notably, the apparent higher safety of the fine-tuned baseline (SG00-FT) in some categories (e.g. CV) is an artifact of survival bias. Its low collision rates are a consequence of poor navigational capability (RC 48.22), see Fig. 6 (right), and high timeout (TO) rates. By failing to progress, the vehicle avoids the complex traffic interactions required for route completion. In contrast, SG10 achieves the highest RC (56.08), demonstrating successful navigation through challenging scenarios, while reducing vehicle collisions by 58% relative to the pretrained checkpoint, providing the best overall balance between safety and route adherence.
Table 4: Explicit relational conditioning. SG11 using Actor-Only graphs with V2 and V3 prompt templates.
Serialization LMDrive BEVDriver